The gap width is 1.400 cm when the temperature is 30.0°C. Determine the gap width when the temperature is 210°C.
Answer #1

delta T = 210 - 30 = 180 0C

L = L0 (1 + alpha * delta T)

= (1.400 cm) (1 + (11 * 10-6 * 180))

= 1.400 * 1.00198

gap width = 1.403 cm